OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (6, 2) and (4, 8)?

ganeshie8 (ganeshie8):

As a start : find the slope from given points

OpenStudy (anonymous):

@ganeshie8 I found the slope and its like this.

ganeshie8 (ganeshie8):

yes, whats the slope ?

ganeshie8 (ganeshie8):

use the slope formula http://www.algebra-class.com/image-files/slope-formula-2.gif

ganeshie8 (ganeshie8):

(6, 2) and (4, 8) x1 y1 x2 y2

ganeshie8 (ganeshie8):

slope = (y2 - y1)/(x2 - x1) = (8 - 2)/(4 - 6) = ?

OpenStudy (anonymous):

Ahh, (2,-6) right?

OpenStudy (anonymous):

2/-6.

ganeshie8 (ganeshie8):

slope is a single number...

ganeshie8 (ganeshie8):

slope = (y2 - y1)/(x2 - x1) = (8 - 2)/(4 - 6) = 6/-2 = -3

ganeshie8 (ganeshie8):

slope = -3

OpenStudy (anonymous):

Yeah, but what about this? y = ? x + ? fill in the ?

OpenStudy (anonymous):

Its from this question.

ganeshie8 (ganeshie8):

slope intercept form of line is : \[\large y = \color{Red}{m}x + \color{Red}{b}\]

ganeshie8 (ganeshie8):

\(\large \color{red}{m}\) : slope \(\large \color{red}{b}\) : y intercept

ganeshie8 (ganeshie8):

you know slope already, right ?

OpenStudy (anonymous):

Yes?

ganeshie8 (ganeshie8):

you just need to find the y intercept

ganeshie8 (ganeshie8):

\[\large y = \color{Red}{-3}x + \color{Red}{b}\]

ganeshie8 (ganeshie8):

plugin any of the given points into above equation and solve `b`

ganeshie8 (ganeshie8):

maybe plugin the first point (6, 2) : \[\large 2 = \color{Red}{-3}(6) + \color{Red}{b}\] solve `b`

OpenStudy (anonymous):

Well, i just dont understand what is b.

ganeshie8 (ganeshie8):

you need to solve the above equation for `b`

ganeshie8 (ganeshie8):

\[\large 2 = \color{Red}{-3}(6) + \color{Red}{b}\] \[\large 2 = -18 + \color{Red}{b}\] \[\large 20 = \color{Red}{b}\]

OpenStudy (anonymous):

I still got more anyone care to look?

ganeshie8 (ganeshie8):

everyone here is for helping only, however you need to try a bit and ask only the part that doesn't make sense..

ganeshie8 (ganeshie8):

i have noticed that you're stuck at solving `2 = -3(6) + b` haven't you dealt with solving equations like that before ?

OpenStudy (anonymous):

No not really.

ganeshie8 (ganeshie8):

you need to know how to solve such equations, they show up everywhere...